Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
pinto, spinicelli - 18:30 Monday 04 December 2023 (62635) Print this report
ISC noise budget and OS monitor/SRCL_SET servo improvement - part 1

This morning we worked on the improvement of the accuracy of the signal used for the OS monitor, SRCL_SET servo. In the first part of the work, we changed online the DARM_LF line (close to the DARM UGF line) and the DARM_LLF line. The idea is to use the latter for the OS monitoring instead of DARM_UGF_line and the DARM_UGF line for the DCP monitoring instead of the DARM_LF. 
Then, we prepared the code ACL for the new loop. In particular:

  • in DARM_ALGO:  moved the DARM_LLF from 34.7Hz to 33.7 Hz (line 26-27)
  • in LSC_ACL_MONI: added demodulation lines for the DARM_LLF (line 28, lines 509-512)
  • in LSC_ACL_MONI: changed the DARM_DCP_LF from DARM_LF to DARM_LINE (commented for the time beeing)
  • in LSC_ACL: added the section for the OS monitoring using the DARM_LLF (lines 697-717). Commented for now; to be used commenting lines 676-696.

In order to use the new loops, we need to uncomment the previous lines and restart LSC_ACL/LSC_ACL_MONI and DARM_ALGO servers. However, since this will also require the tuning of the SR_AA (both in CARM_NULL and LN) plus the calibration of the OS/DCP signals, we decided to postpone this action to not perturbe the following shift. 

After we performed the online test, and performed the modification on Acl, we profited of the stable lock to produce a LSC and ASC noise budget.
At 11.17.06 we went to LowNoise 3.

For the angular loops, we modified the shape of the ASC_noise_butter filter which was previoulsy modified for a SIB1 injection campaign. We moved the butterworth with HP and LP frequencies at 5 and 50 Hz respectively, in order to better measure close to the UGFs.

Here are reported the GPS time stamps of the injections:

Clean data: - 13.50.00 UTC +120 sec


- MICH: 11.22.30 UTC +120 sec, ampl 3e-3
- SRCL: 11.29.30 UTC +120 sec, ampl 2e-3
- PRCL: 11.37.45 UTC +120 sec, ampl 600

- PR TY: 11.52.00 UTC +90 sec, ampl 8e-5
- PR TX: 12.02.40 UTC +90 sec, ampl 8e-5
- DIFFp TX: 12.10.30 UTC +90 sec, ampl 3e-5
- DIFFp TY: 12.18.30 UTC +90 sec, ampl 4e-4

Preliminary projections on Hrec are reported in the attached pictures. Note that all the feedforward filters have not been updated yet.

After the measurements, we reversed the modification in Acl, and let Ilaria start the PR Chrocc tuning.

Some time needs to be reallocated to perform the tests of today with the new frequency line, also to have the possibility to compute the several calibration for the monitors.

Images attached to this report
Search Help